OpenStudy (anonymous):

Find an equation of the line through the given points. Write the equation in standard form. Through (7, -5) and (0, -3) Choose one answer. a. 12x + 3y = -9 b. 2x + 7y = -21 c. -2x + 7y = -21 d. -12x - 3y = -9

OpenStudy (anonymous):

Did you change your screen name, or many people use this same picture?

OpenStudy (anonymous):

b. plug in (7, -5) and (0, -3) to each equation

OpenStudy (anonymous):

First find slope, which is: (y2-y1)/(x2-x1) = (-3+5)/(0-7) = -2/7. Then, find y-intercept in the equation y = -2x/7 + b (where b is y-intercept) using one of the points.

OpenStudy (anonymous):

* and if its the solution you'll get the RHS

OpenStudy (anonymous):

Then you can re-arrange into slope intercept. Alternatively, you can derive the equation yourself like GT posted.

OpenStudy (ash2326):

Equation of a line passing through two point (x2, y2) and (x1, y1) is given by \[ (y-y2)/(x-x2)=(y2-y1)/(x2-x1)\] \[ (x2, y2)= (7, -5)\] \[ (x1, y1)= (0, -3)\] \[(y+5)/(x-7)=(-5+3)/(7-0)\] \[(y+5)/(x-7)=-2/7\] \[7y+35=-2x+14\] \[2x+7y=-21\] option B